# From Source Properties to Strong-Field Tests: a multipronged analysis of GW250114 with an effective one-body model for generic orbits

### v1

@@ Abstract
Introduction: With an exceptional network signal-to-noise ratio (SNR) of $\gtrsim 75$, GW250114_082203 (GW250114, hereafter)—the loudest gravitational-wave (GW) signal of the first decade of GW astronomy—offers unprecedented statistical power for precision studies of binary black hole (BBH) mergers and for performing the most stringent GW-based tests of general relativity (GR) to date (LIGOScientific:2025epi; LIGOScientific:2025obp; Kumar:2025jio; Andres-Carcasona:2025bni). Such a high-SNR signal exemplifies the remarkable sensitivity achieved by the ground-based GW detectors and the opportunities that future facilities promise (Cahillane:2022pqm; Capote:2024rmo), particularly in the context of multi-phase tests of GR (Krishnendu:2021fga; Yunes:2025xwp). However, the statistical precision enabled by such signals places stringent demands on all aspects of the analysis. In particular, waveform accuracy becomes critical: even small modelling inaccuracies can bias the inferred source properties or masquerade as apparent deviations from GR (Hu:2022bji; Chandramouli:2024vhw; Chandra:2024dhf; Gupta:2024gun; Dhani:2024jja; LIGOScientific:2025rsn; LIGOScientific:2025cmm). Similarly, neglecting physical effects such as orbital eccentricity or spin-induced precession could in principle bias parameter estimates for certain sources.
@@
The LIGO-Virgo-Kagra Collaboration (LVK) discovery and companion analyses established GW250114 as a quasi-circular, low-spin BBH merger with source-frame component masses $m_{1}=33.6^{+1.2}_{-0.8}\,M_{\odot}$ and $m_{2}=32.2^{+0.8}_{-1.3}\,M_{\odot}$, and spins $\chi_{1},\chi_{2}\leq 0.24,0.26$ (90% credible intervals) (LIGOScientific:2025epi; LIGOScientific:2025obp). These estimates were shown to be robust across four state-of-the-art non-eccentric BBH waveform models (Varma:2019csw; Pompili:2023tna; Khalil:2023kep; vandeMeent:2023ols; Ramos-Buades:2023ehm; Pratten:2020ceb; Colleoni:2024knd). Using eccentric aligned-spin models, the LVK further constrained the orbital eccentricity to $e\leq 0.03$ at 13.33 Hz (Gamboa:2024hli; Nagar:2024dzj), ascertaining the signal’s non-eccentric origin. Their post-inspiral analyses reported positive evidence for the $(4,\pm 4)$ subdominant mode ($\log_{10}\mathcal{B}=0.54^{+0.18}_{-0.18}$) and demonstrated that the post-peak emission is fully consistent with the Kerr $220$ quasi-normal mode (QNM) and its first overtone (Carullo:2019flw; Isi:2019aib). Residual analyses using best-fit waveforms and inspiral-merger-ringdown (IMR) consistency tests demonstrated agreement with GR, with the latter yielding constraints comparable to those obtained by hierarchically combining events in GWTC-4 (LIGOScientific:2025tgr) and proving that the event satisfied Hawking’s area increase law (Hawking:1971tu). Finally, parameterised plunge-merger-ringdown tests within the pSEOBNR framework found both the $220$ and $440$ QNMs consistent with GR (Brito:2018rfr; Ghosh:2021mrv; Maggio:2022hre; Pompili:2025cdc), while inspiral-phase tests based on post-Newtonian (PN) deviation parameters (Agathos:2013upa) and principal-component analyses (Saleem:2021nsb) reinforced that GW250114 is a textbook example of a GR-consistent BBH signal.
@@
Here we extend the LVK analyses by studying GW250114 with TEOBResumS-Dalí (Gamba:2024cvy; Nagar:2024oyk; Albanesi:2025txj), a state-of-the-art Effective-One-Body (EOB) model capable of describing generic BBH orbits. We quantify the statistical evidence of the various physical effects modelled (eccentricity, spin precession or a combination of the two), demonstrate decisively the presence of the $(4,\pm 4)$ multipoles, and perform three independent tests of GR complementary to those presented by the LVK.
@@
TEOBResumS-Dalí semi-analytically models the radiation multipoles or GW modes $h_{\ell m}(t)$ of a BBH system through a two-stage approach within the EOB framework (Buonanno:1998gg; Buonanno:2000ef)  [^1] Our model version includes the $(\ell,m)=(2,\pm 2),(2,\pm 1),(3,\pm 3),(4,\pm 4)$ multipoles. The inspiral-plunge phase of the modes, $h_{\ell m}^{\mathrm{insp-plunge}}(t)$, is computed from the EOB dynamics via factorized and resummed PN expansions and numerical relativity (NR) calibrations, which capture the non-quasicircular effects during the final plunge (Damour:2008gu; Damour:2007xr; Nagar:2024oyk). This is then smoothly connected to a phenomenologically constructed merger-ringdown template (Damour:2014yha; DelPozzo:2016kmd) $h_{\ell m}^{\mathrm{merger-RD}}(t)$ at mode-dependent matching times $t_{\ell m}^{\mathrm{match}}$ (Nagar:2020pcj):
@@
where the latter combines inputs from black hole (BH) perturbation theory and NR simulations. TEOBResumS-Dalí differs from other EOB-based models (e.g. those presented in Refs. (Pompili:2023tna; Khalil:2023kep; vandeMeent:2023ols; Ramos-Buades:2023ehm; Gamboa:2024imd; Gamboa:2024hli)) in the choices made in its construction, from the resummations adopted to – most prominently – the treatment of non-quasicircular corrections and the merger-ringdown attachment procedure. For further details, we refer the reader to Gamboa:2024hli; Nagar:2024oyk
To probe potential beyond-GR effects within the parameterised IMR framework, we employ two complementary approaches: (1) introduce fractional mode–amplitude deviation parameters, $\delta h_{\ell m}$, and constrain them directly from the data (Puecher:2022sfm; Gupta:2025xxx); (2) apply pTEOBResumS-Dalí (Chiaramello:2025bhi), which perturbs selected NR-calibrated parameters in TEOBResumS-Dalí’s dynamics and waveform generation. This second method is conceptually similar to the parameterised-deviation scheme in the pSEOBNR framework (Brito:2018rfr; Ghosh:2021mrv; Maggio:2022hre; Pompili:2025cdc), used in LVK analyses of GW250114 and earlier events (LIGOScientific:2020tif; LIGOScientific:2021sio) to constrain deviations in QNM frequencies and test the IMR-faithfulness of the ringdown within GR.
In our analysis, pTEOBResumS-Dalí focuses on three key parameters: $\delta M_{f}$ and $\delta\chi_{f}$, which modify the NR-informed remnant mass and spin, and $\delta A_{22}^{\mathrm{peak}}$, which alters the peak amplitude of the dominant $(2,\pm 2)$ mode. Changes in $\delta M_{f}$ and $\delta\chi_{f}$ affect the complex QNM frequencies of $h_{\ell m}^{\mathrm{merger-RD}}(t)$, leading to measurable shifts in the ringdown phase and damping times, while $\delta A_{22}^{\mathrm{peak}}$ perturbs the waveform amplitude near merger. Full details of the pTEOBResumS-Dalí implementation are given in Chiaramello:2025bhi.
@@
Astrophysical nature of GW250114’s source: The LVK discovery and companion analyses established that GW250114 originated from a quasi-circular, low-spin BBH with negligible eccentricity (LIGOScientific:2025epi; LIGOScientific:2025obp). Here we re-examine this astrophysical picture using an extended set of waveform hypotheses, to independently validate and quantify the LVK conclusions, and assess the robustness of the source characterisation across a broader modelling space. To do this, we analyse 8 s of publicly available Advanced LIGO data around GW250114 using four different BBH source hypothesis (see the Supplementary Material (SM) for details) (LIGOScientific:2025snk): (1) non-eccentric, aligned-spin; (2) non-eccentric, precessing-spin; (3) eccentric, aligned-spin; (4) eccentric, precessing-spin.
Across all cases, the data favour a quasi-spherical, i.e., a non-eccentric but precessing configuration, consistent with the LVK findings, which didn’t report statistical evidence. Allowing for spin precession yields a slightly higher likelihood; removing precession reduces the maximum likelihood by $\Delta\max\ln\mathcal{L}\simeq-0.6$ and the Bayesian evidence by $\Delta\log_{10}\mathcal{B}=-0.12\pm 0.11$. Eccentric models—whether aligned-spin or precessing—perform worse, with $\Delta\max\ln\mathcal{L}\simeq-1$ and $\Delta\log_{10}\mathcal{B}\lesssim-1.55\pm 0.12$, and are thus more strongly disfavoured. Our eccentricity constraints also closely match those obtained by the LVK (LIGOScientific:2025epi; LIGOScientific:2025obp). Figure 1 shows the joint posterior on orbital eccentricity $e$ and mean anomaly $\zeta$ obtained using uniform priors $e\in(0,0.4)$ and $\zeta\in(0,2\pi)$. We obtain $e=0.01^{+0.01}_{-0.01},$. Together with the Bayes factors quoted above, these results strengthen LVK’s claim that GW250114 is non-eccentric.
@@
Under the preferred (quasi-spherical) hypothesis, we recover source parameters consistent with the LVK’s estimates. We infer a source-frame total mass of $M_{T}=65.7^{+0.8}_{-0.9}\,M_{\odot}$ and mass ratio $q=1.03^{+0.06}_{-0.03}$, with low component spins, $\chi_{1}=0.13^{+0.11}_{-0.11}$ and $\chi_{2}=0.11^{+0.16}_{-0.09}$, as shown in Fig. 2. These correspond to individual masses of $m_{1}=33.3^{+0.9}_{-0.7}\,M_{\odot}$ and $m_{2}=32.2^{+0.7}_{-1.0}\,M_{\odot}$. Based on these estimates and data-driven GW-based population inference results (KAGRA:2021duu), we predict that the progenitors are likely first-generation (1G) BHs with $\geq 96\%$ certainty (see the SM for details).
@@
From the posteriors and NR-informed fits (Hofmann:2016yih; Jimenez-Forteza:2016oae), we predict that the merger remnant has a final mass $M_{f}=63^{+1}_{-1}\,M_{\odot}\$and spin $\chi_{f}=0.67^{+0.01}_{-0.01}$. Using the surrogate fits of Varma:2019csw; Varma:2020nbm, whose consistency with TEOBResumS-Dalí we have explicitly verified across the posterior-supported region (see the SM for details), we estimate that the remnant’s recoil speed is $v_{\mathrm{kick}}=141_{-117}^{+324}\,\mathrm{km/s}$. The remnant’s mass places it within the predicted pair-instability gap of $60-130\,M_{\odot}$ (Woosley:2021xba; Farmer:2019jed), and depending on the host environment, it may have either been ejected from or retained by it (Gerosa:2021mno).
@@
To ascertain the presence of the $(4,\pm 4)$ modes, we compare the full–multipolar TEOBResumS-Dalí model with an otherwise identical version excluding these harmonics. The run including $(4,\pm 4)$ yields a higher maximum log-likelihood, with $\Delta\ln\mathcal{L}_{\max}\simeq 11.5$, and a modest Bayes factor $\log_{10}\mathcal{B}\simeq 0.66$, with the maximum SNR gain ($\approx 4.79$) being comparable to the “optimal” SNR upper limit obtained for these modes in LIGOScientific:2025epi using the formalism of Mills:2020thr. The modest Bayes factor, despite the substantial log-likelihood increase, reflects that the higher-harmonic improvement is restricted to a narrow region of parameter space and is therefore penalised by model complexity.
@@
We strengthen this finding by identifying when this multipole makes the most significant contribution to the signal. To this end, we use a time-domain likelihood (see SM), progressively excluding data before a start time $\Delta t^{\mathrm{geo}}_{\mathrm{start}}$ (Chandra:2025ipu), defined relative to the signal peak in units of the remnant black-hole mass. At each time step, we compare the TEOBResumS-Dalí model in its quasi-spherical configuration, including both the quadrupole and $(4,\pm 4)$ multipole, against its quadrupole-only analogue. We fix the sky location and peak time $t_{\mathrm{peak}}^{\mathrm{geo}}$ to their maximum-likelihood values for this analysis. As shown in Fig. 4, the $(4,\pm 4)$ multipole has a detectable imprint even in the post-peak portion of the signal, up to $\Delta t^{\mathrm{geo}}_{\mathrm{start}}/M=5$, beyond which the maximum matched-filter SNR relative to a quadrupole-only model drops below 3.
Moreover, as the plot shows, restricting to late-time data yields $\log_{10}\mathcal{B}\sim 1-4$ against the quadrupole-only hypothesis, larger than the one quoted from the full inspiral-merger-ringdown analysis. This is because in a full signal analysis, the earlier inspiral dilutes the relative information carried by the $(4,\pm 4)$ mode. These values are also larger than those obtained from the KerrPostmerger analysis of LIGOScientific:2025epi; we attribute this difference primarily to prior choices and model features.  [^3] KerrPostmerger samples over each mode’s phase at merger independently, whereas in our framework these phases are fixed by the inspiral–plunge dynamics (Gennari:2023gmx). Additionally, KerrPostmerger employs a broader mass prior.444pSEOBNR analyses indirectly suggest such a contribution by constraining deviations in the 440 QNM frequencies, and by comparing the results with simulated signals in which the $(4,\pm 4)$ multipole is removed, although—as quoted in the paper—“the analysis cannot by itself establish whether the 440 QNM is present in the data”.
Figure 5: Residual distributions for Hanford (left) and Livingston (right) detectors after subtracting the whitened TEOBResumS-Dalí posterior waveforms from whitened strain data. Individual curves show posterior samples, colored by their Jensen–Shannon divergence from the expected white noise distribution (black dashed). We find that the residuals are statistically indistinguishable from noise ($\text{JSD}=0.088\pm 0.003$, $p>0.05$ for all samples), thereby indicating that the data is well described by GR-consistent BBH waveforms.
@@
Signal’s consistency with GR: We begin by assessing the consistency of the data with the GR prediction using a modified residual analysis. The LVK companion paper performed this test by subtracting the maximum-likelihood waveform from the data and showing that the residual network SNR is compatible with the noise background with a Kolmogorov–Smirnov p-value of 0.34. Here we complement that study by performing a slightly different test: instead of subtracting a single waveform, we subtract the full ensemble of posterior-supported quasi-spherical TEOBResumS-Dalí waveforms from the data. This approach avoids relying on the single maximum-likelihood waveform returned by an off-the-shelf stochastic sampler, which is not optimised to locate the true maximum log-likelihood point in high-dimensional parameter space (Speagle:2019ivv).
@@
Figure 5 shows the resulting whitened residuals, obtained by subtracting the whitened posterior waveforms from the whitened strain and comparing the distribution with a zero-mean univariate Gaussian. Across the posterior ensemble, the residuals are statistically indistinguishable from the expected coloured Gaussian noise. We measure Jensen–Shannon divergences of $0.088\pm 0.003$ and Kolmogorov–Smirnov $p$-values of $0.23\pm 0.24$, all satisfying $\mathrm{JSD}<0.09$ and $p>0.05$. These show that the general-relativistic TEOBResumS-Dalí waveform model fully captures the gravitational-wave signal and the remaining fluctuations are consistent with detector noise.
@@
Figure 7: Posterior distributions of the fractional $(4,4)$-mode amplitude deviation $\delta h_{44}$. The blue curve shows the observed event, the pink curve a NRSur7dq4 synthetic signal recovered with TEOBResumS-Dalí in zero noise, and the black curve a matched $TEOBResumS-Dalí$ recovery. In the matched case, the $(4,4)$ mode morphology is consistent, yielding a single-peaked posterior near the GR value. For the mismatched synthetic signal and the event, differences in $(4,4)$ mode modelling break the exact $\pi$-phase correspondence, shifting the degenerate solutions to $x$ and x2x-2 ($x\neq 0$) and producing the observed bimodality.
@@
Assuming that the dominant quadrupole $(\ell,m)=(2,\pm 2)$ mode agree with GR, we next test the subdominant $(4,\pm 4)$ multipole under a quasi-circular signal hypothesis.  [^5] As shown earlier, precessing and non-precessing configurations are statistically indistinguishable for this event, supporting the use of the simpler non-precessing analysis. Also, precession leads to significant power mixing between modes of fixed $\ell$, indicating one needs to add these deviations in the co-precessing frame, which we don’t. Following Puecher:2022sfm; Gupta:2025xxx, we introduce a time-independent fractional amplitude deviation $\delta h_{44}$ via $|h_{44}|\ \rightarrow\ |h_{44}|\,(1+\delta h_{44})$, and constrain it directly from the data while enforcing $h_{\ell,-m}=(-1)^{\ell}h_{\ell m}^{\ast}$.
Figure 7 shows the $\delta h_{44}$ posteriors, which are bi-modal. This bimodality reflects an amplitude-phase degeneracy, while the displacement from the GR value $\delta h_{44}=0$ is likely attributable to waveform systematics rather than a true GR violation. To test the latter, we perform two zero-noise simulations with a GW250114-like signal: (i) matched $TEOBResumS-Dalí$ recovery, and (ii) unmatched NRSur7dq4TEOBResumS-Dalí recovery. For an even-$m$ mode-dominated signal like GW250114, the matched analysis exhibits an azimuthal degeneracy $\phi\rightarrow\phi+\pi$, leaving the even-$m$ content unchanged and producing two $\pi$-separated azimuthal islands, both corresponding to $\delta h_{44}\simeq 0$. As a result, the marginal posterior is peaked near the GR value. The unmatched case breaks the “exact” $\pi$-phase correspondence, shifting $\phi$ to absorb part of the waveform difference, producing two displaced $\delta h_{44}$ solutions—mirroring the event and proving the waveform systematic conjecture.
Folded-phase posteriors further support this interpretation: in the matched case, $(\delta h_{44},\phi)$ collapses into two narrow $\pi$-separated bands at $\delta h_{44}\simeq 0$, while in the mismatched and event cases the bands are shifted and broadened. Reflecting the $\delta h_{44}$ posteriors about 1-1, we find that the GR-predicted value lies at the 5.5% credible level of the transformed distribution, with a posterior density of $2.3\times 10^{-1}$ at $\delta h_{44}=0$, within the high-probability region. To our knowledge, this is one of the first attempts to directly test the subdominant $(4,\pm 4)$ multipole’s amplitude consistency with GR in a BBH signal.

### v2

@@ Abstract
Introduction: With an exceptional network signal-to-noise ratio (SNR) of $\gtrsim 75$, GW250114_082203 (GW250114, hereafter)—the loudest gravitational-wave (GW) signal of the first decade of GW astronomy—offers unprecedented statistical power for precision studies of binary black hole (BBH) mergers and for performing the most stringent GW-based tests of general relativity (GR) to date (Abac and others, 2025c, a; Kumar et al., 2025; Andrés-Carcasona and Caneva Santoro, 2025). Such a high-SNR signal exemplifies the remarkable sensitivity achieved by the ground-based GW detectors and the opportunities that future facilities promise (Cahillane and Mansell, 2022; Capote and others, 2025), particularly in the context of multi-phase tests of GR (Krishnendu and Ohme, 2021; Yunes et al., 2025). However, the statistical precision enabled by such signals places stringent demands on all aspects of the analysis. In particular, waveform accuracy becomes critical: even small modelling inaccuracies can bias the inferred source properties or masquerade as apparent deviations from GR (Q. Hu and J. Veitch (2023); R. S. Chandramouli, K. Prokup, E. Berti, and N. Yunes (2025); K. Chandra (2025); A. Gupta et al. (2024); A. Dhani, S. H. Völkel, A. Buonanno, H. Estelles, J. Gair, H. P. Pfeiffer, L. Pompili, and A. Toubiana (2025); A. G. Abac et al. (2025b); 47). Similarly, neglecting physical effects such as orbital eccentricity or spin-induced precession could in principle bias parameter estimates for certain sources.
@@
The LIGO-Virgo-Kagra Collaboration (LVK) discovery and companion analyses established GW250114 as a quasi-circular, low-spin BBH merger with source-frame component masses $m_{1}=33.6^{+1.2}_{-0.8}\,M_{\odot}$ and $m_{2}=32.2^{+0.8}_{-1.3}\,M_{\odot}$, and spins $\chi_{1},\chi_{2}\leq 0.24,0.26$ (90% credible intervals) (Abac and others, 2025c, a). These estimates were shown to be robust across four state-of-the-art non-eccentric BBH waveform models (Varma et al., 2019; Pompili and others, 2023; Khalil et al., 2023; van de Meent et al., 2023; Ramos-Buades et al., 2023; Pratten and others, 2021; Colleoni et al., 2025). Using eccentric aligned-spin models, the LVK further constrained the orbital eccentricity to $e\leq 0.03$ at 13.33 Hz (Gamboa and others, 2025; Nagar et al., 2024), ascertaining the signal’s non-eccentric origin. Their post-inspiral analyses reported positive evidence for the $(4,\pm 4)$ subdominant mode ($\log_{10}\mathcal{B}=0.54^{+0.18}_{-0.18}$) and demonstrated that the post-peak emission is fully consistent with the Kerr $220$ quasi-normal mode (QNM) and its first overtone (Carullo et al., 2019; Isi et al., 2019). Residual analyses using best-fit waveforms and inspiral-merger-ringdown (IMR) consistency tests demonstrated agreement with GR, with the latter yielding constraints comparable to those obtained by hierarchically combining events in GWTC-4 (Abac and others, ) and proving that the event satisfied Hawking’s area increase law (Hawking, 1971). Finally, parameterised plunge-merger-ringdown tests within the pSEOBNR framework found both the $220$ and $440$ QNMs consistent with GR (Brito et al., 2018; Ghosh et al., 2021; Maggio et al., 2023; Pompili et al., 2025), while inspiral-phase tests based on post-Newtonian (PN) deviation parameters (Agathos et al., 2014) and principal-component analyses (Saleem et al., 2022) reinforced that GW250114 is a textbook example of a GR-consistent BBH signal.
@@
Here we extend the LVK analyses by studying GW250114 with TEOBResumS-Dalí (Gamba et al., 2024; Nagar et al., 2025; Albanesi et al., 2025), a state-of-the-art Effective-One-Body (EOB) model capable of describing generic BBH orbits. We quantify the statistical evidence of the various physical effects modelled (eccentricity, spin precession or a combination of the two), demonstrate decisively the presence of the $(4,\pm 4)$ multipoles, and perform three independent tests of GR complementary to those presented by the LVK.
@@
TEOBResumS-Dalí semi-analytically models the radiation multipoles or GW modes $h_{\ell m}(t)$ of a BBH system through a two-stage approach within the EOB framework (Buonanno and Damour, 1999, 2000)  [^1] Our model version includes the $(\ell,m)=(2,\pm 2),(2,\pm 1),(3,\pm 3),(4,\pm 4)$ multipoles. The inspiral-plunge phase of the modes, $h_{\ell m}^{\mathrm{insp-plunge}}(t)$, is computed from the EOB dynamics via factorized and resummed PN expansions and numerical relativity (NR) calibrations, which capture the non-quasicircular effects during the final plunge (Damour et al., 2009; Damour and Nagar, 2007; Nagar et al., 2025). This is then smoothly connected to a phenomenologically constructed merger-ringdown template (Damour and Nagar, 2014; Del Pozzo and Nagar, 2017) $h_{\ell m}^{\mathrm{merger-RD}}(t)$ at mode-dependent matching times $t_{\ell m}^{\mathrm{match}}$ (Nagar et al., 2020):
@@
where the latter combines inputs from black hole (BH) perturbation theory and NR simulations. TEOBResumS-Dalí differs from other EOB-based models (e.g. those presented in Refs. (Pompili and others, 2023; Khalil et al., 2023; van de Meent et al., 2023; Ramos-Buades et al., 2023; Gamboa et al., 2025; Gamboa and others, 2025)) in the choices made in its construction, from the resummations adopted to – most prominently – the treatment of non-quasicircular corrections and the merger-ringdown attachment procedure. For further details, we refer the reader to Gamboa and others (2025); Nagar et al. (2025)
To probe potential beyond-GR effects within the parameterised IMR framework, we employ two complementary approaches: (1) introduce fractional mode–amplitude deviation parameters, $\delta h_{\ell m}$, and constrain them directly from the data (Puecher et al., 2022; Gupta and others, 2025); (2) apply pTEOBResumS-Dalí (Chiaramello et al., 2025), which perturbs selected NR-calibrated parameters in TEOBResumS-Dalí’s dynamics and waveform generation. This second method is conceptually similar to the parameterised-deviation scheme in the pSEOBNR framework (Brito et al., 2018; Ghosh et al., 2021; Maggio et al., 2023; Pompili et al., 2025), used in LVK analyses of GW250114 and earlier events (Abbott and others, 2021, 2025) to constrain deviations in QNM frequencies and test the IMR-faithfulness of the ringdown within GR.
In our analysis, pTEOBResumS-Dalí focuses on three key parameters: $\delta M_{f}$ and $\delta\chi_{f}$, which modify the NR-informed remnant mass and spin, and $\delta A_{22}^{\mathrm{peak}}$, which alters the peak amplitude of the dominant $(2,\pm 2)$ mode. Changes in $\delta M_{f}$ and $\delta\chi_{f}$ affect the complex QNM frequencies of $h_{\ell m}^{\mathrm{merger-RD}}(t)$, leading to measurable shifts in the ringdown phase and damping times, while $\delta A_{22}^{\mathrm{peak}}$ perturbs the waveform amplitude near merger. Full details of the pTEOBResumS-Dalí implementation are given in Chiaramello et al. (2025).
@@
Astrophysical nature of GW250114’s source: The LVK discovery and companion analyses established that GW250114 originated from a quasi-circular, low-spin BBH with negligible eccentricity (Abac and others, 2025c, a). Here we re-examine this astrophysical picture using an extended set of waveform hypotheses, to independently validate and quantify the LVK conclusions, and assess the robustness of the source characterisation across a broader modelling space. To do this, we analyse 8 s of publicly available Advanced LIGO data around GW250114 using four different BBH source hypothesis (see the Supplementary Material (SM) for details) (Abac and others, 2025d): (1) non-eccentric, aligned-spin; (2) non-eccentric, precessing-spin; (3) eccentric, aligned-spin; (4) eccentric, precessing-spin.
Across all cases, the data favour a quasi-spherical, i.e., a non-eccentric but precessing configuration, consistent with the LVK findings, which didn’t report statistical evidence. Allowing for spin precession yields a slightly higher likelihood; removing precession reduces the maximum likelihood by $\Delta\max\ln\mathcal{L}\simeq-0.6$ and the Bayesian evidence by $\Delta\log_{10}\mathcal{B}=-0.12\pm 0.11$. Eccentric models—whether aligned-spin or precessing—perform worse, with $\Delta\max\ln\mathcal{L}\simeq-1$ and $\Delta\log_{10}\mathcal{B}\lesssim-1.55\pm 0.12$, and are thus more strongly disfavoured. Our eccentricity constraints also closely match those obtained by the LVK (Abac and others, 2025c, a). Figure 1 shows the joint posterior on orbital eccentricity $e$ and mean anomaly $\zeta$ obtained using uniform priors $e\in(0,0.4)$ and $\zeta\in(0,2\pi)$. We obtain $e=0.01^{+0.01}_{-0.01}$. Together with the Bayes factors quoted above, these results strengthen LVK’s claim that GW250114 is non-eccentric.
@@
Under the preferred (quasi-spherical) hypothesis, we recover source parameters consistent with the LVK’s estimates. We infer a source-frame total mass of $M_{T}=65.7^{+0.8}_{-0.9}\,M_{\odot}$ and mass ratio $q=1.03^{+0.06}_{-0.03}$, with low component spins, $\chi_{1}=0.13^{+0.11}_{-0.11}$ and $\chi_{2}=0.11^{+0.16}_{-0.09}$, as shown in Fig. 2. These correspond to individual masses of $m_{1}=33.3^{+0.9}_{-0.7}\,M_{\odot}$ and $m_{2}=32.2^{+0.7}_{-1.0}\,M_{\odot}$. Based on these estimates and data-driven GW-based population inference results (Abbott and others, 2023), we predict that the progenitors are likely first-generation (1G) BHs with $\geq 96\%$ certainty (see the SM for details).
@@
From the posteriors and NR-informed fits (Hofmann et al., 2016; Jiménez-Forteza et al., 2017), we predict that the merger remnant has a final mass $M_{f}=63^{+1}_{-1}\,M_{\odot}\$and spin $\chi_{f}=0.67^{+0.01}_{-0.01}$. Using the surrogate fits of Varma et al. (2019, 2020), whose consistency with TEOBResumS-Dalí we have explicitly verified across the posterior-supported region (see the SM for details), we estimate that the remnant’s recoil speed is $v_{\mathrm{kick}}=141_{-117}^{+324}\,\mathrm{km/s}$. The remnant’s mass places it within the predicted pair-instability gap of $60-130\,M_{\odot}$ (Woosley and Heger, 2021; Farmer et al., 2019), and depending on the host environment, it may have either been ejected from or retained by it (Gerosa and Fishbach, 2021).
@@
To ascertain the presence of the $(4,\pm 4)$ modes, we compare the full–multipolar TEOBResumS-Dalí model with an otherwise identical version excluding these harmonics. The run including $(4,\pm 4)$ yields a higher maximum log-likelihood, with $\Delta\ln\mathcal{L}_{\max}\simeq 11.5$, and a modest Bayes factor $\log_{10}\mathcal{B}\simeq 0.66$, with the maximum SNR gain ($\approx 4.79$) being comparable to the “optimal” SNR upper limit obtained for these modes in Abac and others (2025c) using the formalism of Mills and Fairhurst (2021). The modest Bayes factor, despite the substantial log-likelihood increase, reflects that the higher-harmonic improvement is restricted to a narrow region of parameter space and is therefore penalised by model complexity.
@@
We strengthen this finding by identifying when this multipole makes the most significant contribution to the signal. To this end, we use a time-domain likelihood (see SM), progressively excluding data before a start time $\Delta t^{\mathrm{geo}}_{\mathrm{start}}$ (Chandra and Calderón Bustillo, 2025), defined relative to the signal peak in units of the remnant black-hole mass. At each time step, we compare the TEOBResumS-Dalí model in its quasi-spherical configuration, including both the quadrupole and $(4,\pm 4)$ multipole, against its quadrupole-only analogue. We fix the sky location and peak time $t_{\mathrm{peak}}^{\mathrm{geo}}$ to their maximum-likelihood values for this analysis. As shown in Fig. 4, the $(4,\pm 4)$ multipole has a detectable imprint even in the post-peak portion of the signal, up to $\Delta t^{\mathrm{geo}}_{\mathrm{start}}/M=5$, beyond which the maximum matched-filter SNR relative to a quadrupole-only model drops below 3.
Moreover, as the plot shows, restricting to late-time data yields $\log_{10}\mathcal{B}\sim 1-4$ against the quadrupole-only hypothesis, larger than the one quoted from the full inspiral-merger-ringdown analysis. This is because in a full signal analysis, the earlier inspiral dilutes the relative information carried by the $(4,\pm 4)$ mode. These values are also larger than those obtained from the KerrPostmerger analysis of Abac and others (2025c); we attribute this difference primarily to prior choices and model features.  [^3] KerrPostmerger samples over each mode’s phase at merger independently, whereas in our framework these phases are fixed by the inspiral–plunge dynamics (Gennari et al., 2024). Additionally, KerrPostmerger employs a broader mass prior.444pSEOBNR analyses indirectly suggest such a contribution by constraining deviations in the 440 QNM frequencies, and by comparing the results with simulated signals in which the $(4,\pm 4)$ multipole is removed, although—as quoted in the paper—“the analysis cannot by itself establish whether the 440 QNM is present in the data”.
Figure 5: Residual distributions for Hanford (left) and Livingston (right) detectors after subtracting the whitened TEOBResumS-Dalí posterior waveforms from whitened strain data. Individual curves show posterior samples, colored by their Jensen–Shannon divergence from the expected white noise distribution (black dashed). We find that the residuals are statistically indistinguishable from noise ($\text{JSD}=0.008\pm 3\times 10^{-4}$, $p>0.05$ for all samples), thereby indicating that the data is well described by GR-consistent BBH waveforms.
@@
Signal’s consistency with GR: We begin by assessing the consistency of the data with the GR prediction using a modified residual analysis. The LVK companion paper performed this test by subtracting the maximum-likelihood waveform from the data and showing that the residual network SNR is compatible with the noise background with a Kolmogorov–Smirnov p-value of 0.34. Here we complement that study by performing a slightly different test: instead of subtracting a single waveform, we subtract the full ensemble of posterior-supported quasi-spherical TEOBResumS-Dalí waveforms from the data. This approach avoids relying on the single maximum-likelihood waveform returned by an off-the-shelf stochastic sampler, which is not optimised to locate the true maximum log-likelihood point in high-dimensional parameter space (Speagle, 2020).
@@
Figure 5 shows the resulting whitened residuals, obtained by subtracting the whitened posterior waveforms from the whitened strain and comparing the distribution with a zero-mean univariate Gaussian. Across the posterior ensemble, the residuals are statistically indistinguishable from the expected coloured Gaussian noise. We measure Jensen–Shannon divergences of $0.088\pm 0.003$ and Kolmogorov-Smirnov $p$-values of $0.23\pm 0.24$, all satisfying $\mathrm{JSD}<8\times 10^{-3}$ and $p>0.05$. These show that the general-relativistic TEOBResumS-Dalí waveform model fully captures the gravitational-wave signal and the remaining fluctuations are consistent with detector noise.
@@
Figure 7: Posterior distributions of the fractional $(4,4)$-mode deviation, $\delta h=(1+\delta A_{44}e^{i\delta\phi_{44}})$. The blue curve shows the observed event, the pink curve a NRSur7dq4 synthetic signal recovered with TEOBResumS-Dalí in zero noise, and the green curve a matched $TEOBResumS-Dalí$ recovery. In all cases, the $(4,4)$ mode deviation is consistent with GR, yielding a posterior peaked near the value of 11 (vertical black line). However, in the event and the mismatched synthetic signal, we observe a slight shift of the posterior amplitude to the right, likely due to waveform systematics.
@@
Assuming that the dominant quadrupole $(\ell,m)=(2,\pm 2)$ mode agrees with GR, we next test the subdominant $(4,\pm 4)$ multipole under a quasi-circular signal hypothesis.  [^5] As shown earlier, precessing and non-precessing configurations are statistically indistinguishable for this event, supporting the use of the simpler non-precessing analysis. Also, precession leads to significant power mixing between modes of fixed $\ell$, indicating one needs to add these deviations in the co-precessing frame, which we don’t. Following and extending Puecher et al. (2022); Gupta and others (2025), we introduce a time-independent fractional mode deviation, parameterised by an amplitude and a phase deviation via $h_{44}\ \rightarrow\ h_{44}\,(1+\delta A_{44}e^{i\delta\phi_{44}})$, and constrain it directly from the data while enforcing $h_{\ell,-m}=(-1)^{\ell}h_{\ell m}^{\ast}$. Figure 7 shows the deviations posteriors, which are consistent with the GR prediction of $(1+\delta A_{44}e^{i\delta\phi_{44}})=1$ at the 90% credible level, though the amplitude posterior shows a long tail towards higher values and a shift of the peak away from the GR expectation. To test the robustness of this finding, we perform two additional zero-noise simulations with a GW250114-like signal: (i) matched $TEOBResumS-Dalí$ recovery, and (ii) unmatched NRSur7dq4TEOBResumS-Dalí recovery. The former yields a unimodal posterior peaked at the GR value, while the latter reproduces the shift in in peak amplitude deviation similar to that observed in the event analysis. This suggests that waveform systematics are responsible for this feature. To our knowledge, this is one of the first attempts to directly test the subdominant $(4,\pm 4)$ multipole’s amplitude consistency with GR in a BBH signal.